OpenStudy (anonymous):

I need help on this calculus word problem: Two ants are at a common point at time t=0, the first ant starts crawling along a straight line at the rate of 4 ft/min. Two minutes later, the second ant starts crawling in a direction perpendicular to that of the first, at a rate of 5 ft/min. How fast is the distance between them changing when the first insect has traveled 12 feet?

OpenStudy (amistre64):

what formula can we use to determine the distance between them?

OpenStudy (anonymous):

Pythagorean theorem

OpenStudy (amistre64):

good, implicit that, what do we get?

OpenStudy (amistre64):

d^2 = x^2 + y^2 2d d' = 2x x' + 2y y' d d' = x x' + y y' d' = x x' + y y' -------- d we are told x' and y', and the information is all given to determine x,y,d

OpenStudy (anonymous):

dx/dt=4the rate of the first ant dy/dt=5the rate of the second ant x=12 the distance the first ant traveled.

OpenStudy (anonymous):

distance = rate*time

OpenStudy (amistre64):

good, and 3 minutes has passed to get to 12 feet

OpenStudy (amistre64):

3 min - 2 min = 1 min @ 5ft/min, y = 5

OpenStudy (anonymous):

For 1st Ant t=3

OpenStudy (anonymous):

for 2nd t=5

OpenStudy (amistre64):

t=3 regardless :)

OpenStudy (amistre64):

if we assume the 2nd ant started at the same time, then it was simply at -10 and didnt get to the crossing yet

OpenStudy (anonymous):

We use the Pythagorean Theorem to find r

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

r=13

OpenStudy (anonymous):

thanks for help

OpenStudy (amistre64):

youre welcome